OpenStudy (anonymous):

Evaluate. (34 + 14) ÷ [42 – (20 + 16)] what is the answer

OpenStudy (anonymous):

First solve the brackets.. What is : 34 + 14 = ?? And what is 20 + 16 = ??

OpenStudy (anonymous):

34+14 is 48 and 20+16 is 36 now what do i do

OpenStudy (anonymous):

48÷6 = 8 ....that's ur answer

OpenStudy (jerudavis):

8